### v1

# Difference Decomposition Networks for Infrared Small Target Detection

Abstract
One of main challenges faced by infrared small target detection (ISTD) is weak targets. To enhance target saliency, recent studies have widely adopted difference-based methods, which highlight the weak target by computing and weighting difference features. However, these methods rely on fixed and predetermined weights to combine these difference features in spatial and temporal domains. These approaches fail to adaptively capture the inter-correlations among various directions or different frames. Moreover, these inter-correlations contain crucial structural information about the target. Consequently, this failure limits the model’s performance. To address this limitation, we propose a Basis Decomposition Module (BDM) to extract the correlation information in different directions. For single-frame ISTD (SISTD), we introduce Spatial Difference Decomposition Module (SD2M) and Spatial Difference Decomposition Downsampling Module (SD3M), all based on the BDM structure, to extract the detailed information in the spatial dimensions. In addition, we use the BDM-based Temporal Difference Decomposition Module (TD2M) to extract the motion information for multi-frame ISTD (MISTD). By incorporating our proposed modules into a stacked U-Net architecture, we introduce two networks: the Spatial Difference Network (SD2Net) for SISTD and the Spatio-Temporal Difference Network (STD2Net) for MISTD. These networks are designed to achieve precise detection of infrared small targets without a large number of parameters. On the SISTD task, SD2Net achieves competitive performance against most established networks with 0.27M parameters. On the MISTD datasets, STD2Net achieves a mIoU of 87.68%, outperforming SD2Net, which achieves a mIoU of 64.97%. Our codes are available: https://github.com/greekinRoma/IRSTD_HC_Platform.
I Introduction
Unlike visible light imaging, infrared sensing maintains reliable performance in adverse weather conditions, offering high detection reliability, concealment, and mobility. As one of the main applications of infrared sensing, Infrared Search and Track (IRST) have demonstrated the strong robustness under varying circumstances, and works as the fundament in several applications [Zhang2024IRPruneDet, Zhu2025PFF, iMoPKL2025, MICPL2025].
Infrared small target detection (ISTD) is the core of IRST systems, and has attracted growing research interests. Nevertheless, the dim targets and complex backgrounds render ISTD significantly more difficult than many other detection tasks [Zhu2024GSTUnet]. Therefore, ISTD is a fundamental and challenging computer vision task, and merits further investigation [dai2025SeRankDet, zhang2025daaf, zhang2025coupled, zhang2025dual].
To address these challenges in ISTD, numerous model-based approaches are proposed for single-frame ISTD (SISTD) [kong2022LogTFNN, cao2021infrared, dai2017reweighted, wang2019miss]. With the recent development of deep-learning techniques, data-driven methods have demonstrated superior performance to model-driven methods [Zhang2025IRmamba, li2025hstnet, li2025mmidnet, zhang2025exploringstatespacemodel]. Up to now, the methods based on deep learning have become the mainstream of ISTD.
In recent years, difference-based methods, like central difference convolutions (CDCs) and difference-frame methods, have gained the increasing attention [Wu2025Salm]. In these modules, the detailed information are highlighted using high-pass kernels in spatial or temporal dimensions[Zhang2021ECA]. However, both CDC-based approaches and other spatial-domain methods face a fundamental limitation in distinguishing small targets from structurally similar clutter without the leverage of motion information.
In order to solve above problems, motion information is introduced into the algorithms. Existing general video object detection (GVOD) methods [Deng2019Object, zhu2017deep, wu2019sequence, zhou2022transvod, wang2022ptseformer] are designed for large targets with apparent inter-frame motion. Nevertheless, the infrared small targets are usually dim and move slowly in infrared streams[Duan2024Triple, zhang2025fsatfusion, zhang2025wife], and the GVOD methods could not highlight the motion information of dot targets. Therefore, GVOD approaches are not suitable for ISTD. This motivates the exploration of new algorithms designed for infrared small targets. Currently, the common technique for obtaining motion information for infrared small targets relies on the difference between key frames and reference frames [Li2025dtum].
As illustrated in Figure 1, difference-based methods all involve two steps: first, they extract difference features by subtracting, and then combine these features using coefficients. These coefficients are learned during training but remain static during inference.
((a))
((b))
Figure 1: The standard CDC and frame-difference operations, both employing predetermined static coefficients to combine difference results.
According to basis decomposition, an input can be represented as a combination of basis features and their dynamic coefficients. These input-specific coefficients are unique per basis and inherently capture correlation information between features. To address the limitation of fixed weights in difference-based methods, we introduce basis decomposition. This enables adaptive feature combination based on the learned correlations, thereby improving performance.
Based on basis decomposition, we construct a unified and lightweight framework for ISTD, named the Basis Decomposition Module (BDM). The framework can enhance task-relevant information, like difference information, in both spatial and temporal domains. These coefficients can be regarded as an inherent property of the input features. Therefore, their extraction does not require any additional parameters.
Building on BDM, we design a series of networks collectively referred to as Difference Decomposition Networks (DDNets). For SISTD, we utilize adaptations of BDM in the spatial domain, specifically the Spatial Difference Decomposition Module (SD2M) and the Spatial Difference Decomposition Downsampling Module (SD3M), to enhance high-frequency target features while suppressing low-frequency backgrounds. Based on two modules, we propose the Spatial Difference Decomposition Network (SD2Net). For MISTD, we further extend SD2Net to Spatiotemporal Difference Decomposition Network (STD2Net) by integrating temporal information through the Temporal Difference Decomposition Module (TD2M), which emphasizes the differences among frames and the features of the current frame in the temporal domain. The proposed networks not only outperform existing methods but also have a smaller model size. The contributions are summarized as below:
  1. 1.
    We introduce the BDM, a foundational component that enhances task-relevant features while suppressing task-irrelevant information through basis decomposition.
  2. 2.
    We develop SD2M and SD3M, derived from BDM, into a modified U-Net, proposing SD2Net for SISTD to strengthen high-frequency target features and suppress low-frequency clutter with few parameters.
  3. 3.
    For detecting moving targets against static backgrounds, we propose the TD2M. By incorporating TD2M, we further build STD2Net, which effectively highlights moving targets and suppresses the clutter.
II Related Work
II-A Single-frame Infrared Small Target Detection
Existing SISTD methods can be categorized into two methodological paradigms: model-driven methods and data-driven methods.
Model-driven methods leverage domain-specific priors and bifurcate into three branches. Filter-based methods [Qin2019Facetkernel] enhance target features and suppress clutter with filters.Although these techniques perform well in simple scenes, they are highly sensitive to window size and easily fail when clutter exhibits similar texture or brightness. Human visual system-based methods [WeiYouLi2016Multiscale, Han2021WLCM] are based on visual characteristics of the visual system.However, these methods rely on handcrafted contrast metrics and struggle to generalize across diverse imaging conditions. Optimization-based methods [gao2013infrared, ZhangPeng2019Infrared] formulate ISTD as an optimization problem to detect the targets by separating them from the background.Despite their interpretability, these models usually require iterative solvers, leading to heavy computation and limited scalability Although these methods are computationally efficient and interpretable, they are usually sensitive to parameter selection and fail under complex cluttered scenes.
Data-driven methods have addressed these issues of SISTD by learning discriminative features from large datasets. Dai et al. have successively proposed the Asymmetric Contextual Modulation (ACM) [DaiWuZhouBarnard2021Asymmetric] and Attentional Local Contrast Network (ALCNet) [dai2021attentional] for ISTD. Dense Nested Attention Network (DNANet) [RenLiHanShu2021DNANet] supports adaptive interactions between feature layers to prevent the loss of targets. Furthermore, UIUNet [WuHongChanussot2023UIU-Net] employs a simple and effective “U-Net in U-Net” framework to enhance both global and local contrast information. Receptive-field and Direction-induced Attention Network (RDIAN) [SunBaiYangBai2023Receptive-Field] utilizes the characteristics of targets to solve the imbalance between targets and background. Infrared Small-target Detection U-Net (ISTDU-Net) [HouZhangTanXiZhengLi2022ISTDU-Net] introduces a fully connected layer in the skip connection to suppress the backgrounds with similar structures from the global receptive field. The Attention-guided Pyramid Context Network (AGPCNet) [ZhangCaoPuPeng2021AGPCNet] employs an attention-guided context block, providing the model with a perspective on both inner and global patches. Spatial-Channel Cross Transformer Network (SCTransNet) [Yuan2024SCtransNet] uses the Transformer to obtain the effective global information.
Nevertheless, these architectures tend to be computationally heavy, and the lack of explicit mechanisms for separating target-relevant and background-redundant features often leads to false alarms under complex thermal clutter. Therefore, how to design a lightweight, interpretable, and generalizable representation mechanism remains an open issue in single-frame ISTD.
II-B Multi-frame Infrared Small Target Detection
Multi-frame frameworks can be categorized by their strategies for extracting temporal information.
Pre-extraction frameworks extract motion features before a SISTD network. Energy Accumulation (IFEA) [Du2022IFEA] and Spatio-temporal Differential Multiscale Attention Network (STDMANet) [Yan2023STDMANet] enhance the energy of targets by computing the difference between the current frame and reference frames before input into networks.
Result-based extraction frameworks extract temporal information based on the results of single-frame neural networks. Direction-code Temporal U-shape Module (DTUM) [Li2025dtum] designs a new algorithm to extract the motion features from the results of SISTD networks. These approaches emphasize explicit motion cues and are straightforward to implement, but their reliance on handcrafted frame-difference operators makes them sensitive to background dynamics and sensor noise, often resulting in unstable motion estimation. The Recurrent Feature Refinement (RFR) framework [ying2025rfr] effectively captures long-term temporal dependencies based on the outputs of SISTD networks. Although such methods exploit temporal refinement after initial detection, they often inherit errors from the single-frame stage, causing accumulated false alarms when the early predictions are unreliable.
Spatiotemporal feature simultaneous extraction frameworks obtain the temporal and spatial information using modules in a single network. Sliced Spatio-temporal Network (SSTNet) [Chen2024SSTNet] utilizes the ConvLSTM node to extract the motion information of the small infrared targets. The Lightweight Asymmetric Spatial Feature Network (LASNet) [Chen2024LASNet] utilizes motion feature extraction and motion-affinity fusion for extracting temporal information. The Triple-domain Strategy (Tridos) [Duan2024Triple] employs frequency-aware enhancement to enhance the detection of dim infrared targets. The Local Motion Aware Transformer (LMAFormer) [huang2024lmaformer] utilizes local motion-aware attention to enhance the weak target motion information in MISTD. These end-to-end architectures can effectively integrate spatial and temporal cues, yet they usually demand heavy computation and high memory usage, limiting their deployment in real-time infrared tracking systems. Furthermore, their focus on global aggregation often overlooks feature disentanglement, making small targets difficult to separate from cluttered backgrounds.
Although these multi-frame networks improve robustness, most rely on heuristic motion aggregation or fusion-based temporal modules. They lack a unified and interpretable mechanism to decouple target-relevant and background-redundant components, leading to limited generalization across different ISTD tasks.
In summary, current SISTD and MISTD methods are largely designed for specific subtasks, lacking theoretical extensibility and efficiency. To overcome these limitations, we introduce a Basis Decomposition Module (BDM) that provides a unified, lightweight foundation for both spatial and temporal feature enhancement.
III Methodology
III-A Preliminaries
((a))
((b))
Figure 2: (a) The procedure of basis decomposition. (b) The illustration of the proposed framework for signal recovery and task-relevant feature enhancement. First, we collect both task-relevant and task-irrelevant basis features into a basis pool according to the certain tasks. Then, the input is then processed in two parallel branches for recovery and enhancement. The upper branch indicates the recovery process, where the input is reconstructed through basis decomposition. The lower branch outlines the enhancement procedure, which selects the task-relevant basis features to reconstruct the input.
A vector $\vec{x}\in\mathbb{R}^{c_{\text{ful}}}$ can always be expressed as a linear combination of $c_{\text{ful}}$ orthonormal basis features.
As shown in the upper part of Figure 2, a vector $\vec{x}\in\mathbb{R}^{c_{\text{ful}}}$ can be recovered by $c_{ful}$ orthonormal vectors:
$c_{ful}=dim(\vec{x}),$ (1)
$s_{i}=\vec{b}_{i}^{\top}\vec{x}\,$, (2)
$\vec{x}=\sum_{i=1}^{c_{\text{ful}}}s_{i}\vec{b}_{i}\,$, (3)
where $s_{i}$ denotes the coefficient of $\vec{x}$ projected onto the basis vector $\vec{b}_{i}$. $\top$} is the transposition operation, and $c_{\text{ful}}$ is the dimension of the input vector.
However, by removing redundant basis vectors, the result is constructed solely from the selected basis elements, thereby enhancing the task-relevant features, as illustrated in the lower portion of Figure 2. By retaining only those vectors that are relevant to the task and removing irrelevant ones, we can strengthen the important features while reducing interference from non-essential information.
Therefore, for a given task, it is not necessary to retain all basis features, and we could select the task-relevant basis vectors to obtain $\vec{x}$ and approximate the input vector $\vec{x}$:
$\vec{y}=\sum_{i=1}^{c_{\text{tas}}}s_{i}\vec{b}_{i}\quad\text{where}\quad s_{i}=\vec{b}_{i}^{\top}\vec{x}\,$, (4)
where $c_{\text{tas}}$ is the number of the task-relevant basis features.
In addition, as shown in Table I, our method does not require the basis vectors to be strictly orthonormal. This relaxation of the orthogonality constraint not only simplifies the computation process but also improve the performance.
When it comes to ISTD dependent on comprehensive and detailed information, we could select the difference information to construct the feature. Consequently, we retain the essential features as basis feature vectors to form $\vec{y}$ and approximate $\vec{x}$.
III-B Basis Decomposition Module
We propose the BDM based on the above discussion, as shown in Figure 3.
Figure 3: Overview of the BDM.
Firstly, we employ the Feature Extraction Module (FEM) and Basis Feature Extraction Module (BFEM) to obtain the original feature $T\in\mathbb{R}^{B\times G\times 1\times F}$ and the basis features $P_{i}\in\mathbb{R}^{B\times G\times 1\times F}$ from $I\in\mathbb{R}^{B\times G\times 1\times F_{inp}}$:
$T$ $=\mathrm{FEM}(I)\,$, (5)
$P_{1},\dots,P_{c}$ $=\mathrm{L2Norm}(\mathrm{BFEM}(I))\,$, (6)
where $B$ is the batch size, $F$ is the feature dimension, and $F_{inp}$ is the input feature dimension. $\mathrm{L2Norm}$ refers to L2 normalization to ensure each basis feature has the same influence on $\vec{y}$.
Then, We concatenate the $\{P_{i}\}$ to obtain $P\in\mathbb{R}^{B\times G\times c\times F}$:
$P=[P_{1};\dots;P_{c}]\,$, (7)
where $c$ refers to the number of input channels, while $\cdot$$\cdot$] is the concatenation operator applied along the certain dimension. $G$ represents the number of feature groups, as we can regard several features across channels as one group.
After that, we utilize the $T$ and $P$ to obtain the tensor $S\in\mathbb{R}^{B\times G\times 1\times c}$ from concatenating decomposition coefficients $S_{i}\in\mathbb{R}^{B\times G\times 1\times 1}$ for different basis features $P_{i}$:
[…]
where $[,]$ concatenates elements along the last dimension.
Finally, the weight $S_{i}$ is assigned to each basis feature $P_{i}$:
$\begin{split}O=S\cdot P=(T\cdot P^{\top})\cdot P=\sum_{i=1}^{c}(T\cdot P_{i}^{\top})\cdot P_{i}\,.\end{split}$ (9)
In short, FEM extracts features containing both salient and redundant information, and BDM approximates the FEM results by linearly combining the basis features from BFEM. This approximation suffers from information loss. Therefore, we must retain the essential information and disregard the redundant information. As the definition of essential information is task-dependent, BFEM requires corresponding adjustments for specific tasks.
In processing two-dimensional feature maps, spatial dimensions (height $H$ and width $W$) are flattened into tokens. For an input $\mathbf{I}\in\mathbb{R}^{B\times C\times H\times W}$, this yields basis features of size $B\times C\times c\times(HW)$, where $HW$ and $C$ map to dimensions $F$ and $G$.
((a))
((b))
Figure 4: Detailed structure of the SD2M and SD3M used in Figures 6 and 8 respectively, are constructed based on the BDM.
Figure 5: Our backbone adopts a three-stage U-Module structure, where each U-Module is an encoder–decoder architecture. The first U-Module extracts coarse information, the second U-Module performs as the main feature-extracted stage to obtain the refine information, and the third U-Module conducts fine-grained enhancement.
Figure 6: The overall architecture of SD2Net, which incorporates the SD2M and SD3M modules.
III-C Spatial Difference Decomposition Network
As illustrated in Figures 4(a) and 4(b), we adjust the BDM for ISTD and introduce two modules, named SD2M and SD3M, to enhance the targets and suppress the clutter. In addition, we add those modules to our backbone, as shown in Figure 5, and propose the SD2Net, as illustrated in 6.
III-C1 Backbone
As illustrated in Figure 5, We design an architecture composed of three U-Modules. Given an input $I$, the first U-Module produces a coarse representation $f_{1}=\mathcal{U}_{1}(I)$. The second U-Module extract features further and works as the main feature extraction layer: $f_{2}=\mathcal{U}_{2}(f_{1})$. The third U-Module works for final result enhancement: $O=\mathcal{U}_{3}(f_{2})$. The overall mapping is
$O=\mathcal{U}_{3}\big(\mathcal{U}_{2}(\mathcal{U}_{1}(I))\big).$
Figure 7: Architecture of the DFEM, which employs 8 convolutional kernels with $n$ dilation rates to extract multi-scale basis features.
Figure 8: The architecture of Inception Pooling.
III-C2 Spatial Difference Decomposition Module
As depicted in Figure 4(a), we employ a $1\times 1$ convolution layer followed by a reshape operation ($\mathrm{Reshape}$) as FEM to extract the original feature $T_{m}\in\mathbb{R}^{B\times C\times 1\times(HW)}$:
$T_{m}=\mathrm{Reshape}(\mathrm{Conv}_{1\times 1}(I))\,$, (10)
where H, W, and C denote the height, width, and channel dimension. Difference Feature Extraction Module (DFEM) uses the difference kernels to extract the high-frequency as basis features ${P^{d_{j}}_{i}}\in\mathbb{R}^{B\times C\times 1\times(HW)}$ (where i and j index the kernel type and the dilation rate). These features are then used to form the tensor $P_{m}\in\mathbb{R}^{B\times C\times(8n)\times(HW)}$:
$P_{m}$ $=\mathrm{L2Norm}(DFEM(I))\,$, (11)
$DFEM(I)$ $=[P^{d_{1}}_{1};\ldots;P^{d_{n}}_{1};\ldots;P^{d_{n}}_{8}]\,$, (12)
$P^{d_{j}}_{i}$ $=\text{Reshape}(\mathrm{Conv}_{K_{i}}^{d_{j}}(I))\,$, (13)
where $\mathrm{Conv}_{K_{i}}^{d_{j}}$ is the difference convolution with kernels $K_{i}$ ($i\in\{1,\dots,8\}$), as shown in Figure 7, and dilation rates $d_{j}$ ($j\in\{1,\dots,n\}$) to extract the high-frequency features as our basis features.
We utilize $P_{m}$ and $T_{m}$ to obtain $O_{m}\in\mathbb{R}^{B\times C\times H\times W}$ according to Equation 9:
$O_{m}=\mathrm{Reshape}((T_{m}\cdot P_{m}^{\top})\cdot P_{m}\,)\,$, (14)
where $O_{m}$ is a combination of difference features. In the process, the background information is suppressed by overlooking the low-frequency basis features.
Figure 9: The structure of STD2Net.
III-C3 Inception Pooling.
To maintain feature diversity at various levels, Inception Pooling splits the feature channels into groups and applies distinct operations to each group, as shown in Figure 8.
First, we split the input $I$ along the channel dimension into three equal-sized tensors $C_{t}\in\mathbb{R}^{B\times(C/3)\times H\times W}$.
$C_{1},C_{2},C_{3}=\mathrm{Split}(I)\,$, (15)
After that, we use MaxPooling $\mathrm{Max}_{\text{2\texttimes 2}}$, AvgPooling $\mathrm{Avg}_{\text{2\texttimes 2}}$, and $\text{SD}^{\mathrm{3}}\text{M}$ to downsample the input and concatenate the results in channel dimension by concatenation ($\mathrm{CAT}$).
$O=\text{CAT}(\mathrm{Max}_{\text{2\texttimes 2}}(C_{1}),\mathrm{Avg}_{\text{2\texttimes 2}}(C_{2}),\text{SD}^{\mathrm{3}}\text{M}(C_{3}))\,$, (16)
III-C4 Spatial Difference Decomposition Downsampling Module.
As illustrated in Figure 4(b), the SD3M shares the similar architecture with SD2M, differing in the original feature $T_{d}\in\mathbb{R}^{B\times C\times 1\times(HW/4)}$ and the basis features $P_{d}\in\mathbb{R}^{B\times C\times 4\times(HW/4)}$, which is computed as follows:
$T_{d}$ $=\mathrm{Reshape}(\mathrm{Conv}_{\mathrm{1\times 1}}(\mathrm{Avg}_{\mathrm{2\times 2}}(I)))\,$, (17)
$P_{d}$ $=\mathrm{L2Norm}(\text{DFED}(I))\,$, (18)
where DFED denotes the Difference Feature Extraction Downsampling, formulated as follows:
$\text{DFED}(I)$ $=[P_{max};P_{D_{1}};P_{D_{2}};P_{D_{3}}])\,$, (19)
$P_{max}$ $=\mathrm{Reshape}(\mathrm{Max}_{\mathrm{2\times 2}}(I))\,$, (20)
$P_{D_{k}}$ $=\mathrm{Reshape}(\mathrm{Conv}^{1}_{D_{k}}(I))\,$, (21)
where $\mathrm{Conv}^{1}_{D_{k}}$ denotes a convolution with stride 2, dilation ratio of 1, and the kernel $D_{k}$, which is defined as: $D_{1}=[1,1;-1,-1],D_{2}=[1,-1;1,-1],D_{3}=[1,-1;-1,1]$.
At last, we obtain the output of $O_{d}\in\mathbb{R}^{B\times C\times H/2\times W/2}$:
$O_{d}=\mathrm{Reshape}((T_{d}\cdot P_{d}^{\top})\cdot P_{d}\,)\,.$ (22)
where $O_{d}$ is a linear sum of high-frequency difference features and MaxPooling results from I.
III-D Spatiotemporal Difference Decomposition Network
Since background clutter exhibits random motion patterns, while targets move consistently. Therefore, we introduce TD2M to motion information and propose STD2Net. The basis features extracted by TD2M consist of both the differences in SD2Net’s results between the current frame and the reference frames, and the current frame’s SD2Net output. The former captures motion information, while the latter provides current state information.
Figure 10: Visual results from varied data-driven methods. Green, red, and yellow boxes indicate detected targets, missed targets, and false alarms.
At first, we extract the input features $I_{k}\in\mathbb{R}^{B\times C\times H\times W}$ (k = 1,…,$N$) for each of the $N$ consecutive frames using SD2Net, and concatenate them in channel dimension to obtain $I\in\mathbb{R}^{B\times(NC)\times H\times W}$ :
$I=\mathrm{CAT}(I_{1},\dots,I_{N})\,$, (23)
where the last frame is the current frame and the preceding N-1 frames are reference frames.
Then, we utilize the TD2M to extract the relationships across these consecutive frames. We obtain original features $T_{t}\in\mathbb{R}^{B\times C\times 1\times(HW)}$:
$T_{t}=\mathrm{Reshape}(\mathrm{Conv}_{1\times 1}(I))\,.$ (24)
After that, we employ the Temporal Feature Extraction Module (TFEM) and the normalization layer to obtain the basis features $P_{t}\in\mathbb{R}^{B\times C\times N\times(HW)}$:
$P_{t}$ $=\mathrm{L2Norm}(\mathrm{TFEM}(I))\,$, (25)
$\mathrm{TFEM}(I)$ $=[P_{D_{1}};..;P_{D_{N-1}};P_{C}]\,$, (26)
$P_{D_{i}}$ $=\mathrm{Reshape}(I_{N}-I_{i})\,$, (27)
$P_{C}$ $=\mathrm{Reshape}(I_{N}),$ (28)
where $P_{D_{i}}$ denotes the difference between the current frame $I_{N}$ and the reference frame $I_{i}$. $P_{C}$ is the state of the current frame.
Later, we approximate $T_{t}$ with a combination of $P_{t}$ and obtain the result $O_{t}\in\mathbb{R}^{B\times C\times H\times W}$:
$O_{t}=\mathrm{Reshape}((T_{t}\cdot P_{t}^{\top})\cdot P_{t})\,.$ (29)
At last, we use a U-Module in Figure 6 to further extract information and a convolution layer to adjust the output channel dimension.
Although the dot-product form resembles attention, BDM differs fundamentally: it performs linear basis reconstruction rather than parametric attention weighting. The coefficients arise from geometric alignment between features and fixed bases, ensuring interpretability and low computational cost.
III-E Loss Function
We employ the soft Intersection over Union (IoU) loss for training our model, which is defined as:
$\mathcal{L}=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (30)
where $g_{i,j}\in[0,1]$ and $p_{i,j}\in[0,1]$ represent the ground truth and predicted probability after Sigmoid at $(i,j)$.
IV Experiments
We present the experiment details, validate modules via ablation studies, and demonstrate the performance of BDM’s adaptations through comparisons between SISTD and MISTD.
IV-A Experimental Details
IV-A1 Experiment Dataset
For the task of the SISTD, we employ two widely recognized public benchmarks: NUDT-SIRST [RenLiHanShu2021DNANet] and IRSTD-1K [Zhang2022ISNet]. To evaluate performance on the more complex task of MISTD, we benchmark our network against existing SOTA methods using the comprehensive NUDT-MIRSDT dataset [Li2025dtum]. To ensure a fair and consistent evaluation across all experiments, we divide each dataset into two parts: 80% of the images are randomly selected to constitute the training set, while the remaining 20% are reserved as the test set for the final performance assessment. All comparative methods were evaluated on the same dataset to ensure a fair comparison.
TABLE I: Impact of orthogonality and non-orthogonality.
[…]
TABLE II: Impact of SD2Net’s channel configs on NUDT-SIRST dataset. We bold the best results and underline the second best.
[…]
TABLE III: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.
[…]
IV-A2 Implementation Details
The loss function for all models is the IoU loss [yu2016unitbox]. All experiments are conducted on Ubuntu 20.04 LTS. We train SISTD models on an NVIDIA GeForce RTX 4080 GPU for 400 epochs, using the Adam optimizer with a learning rate of $5\times 10^{-4}$. The batch size is set to 4. The learning rate is decayed to $5\times 10^{-5}$ at epoch 200 and further reduced to $5\times 10^{-6}$. Our MISTD architecture runs on two 24GB GeForce RTX 3090 GPUs. We train for 20 epochs with the Adam optimizer and an initial learning rate of 0.001. Furthermore, we set the temporal window size $N$ of STD2Net to 5. We run each algorithm three times and use the average outcome as the final result.
IV-A3 Evaluation Metrics
We adopt four evaluation metrics: false alarm rate ($\mathrm{P_{a}}$), probability of detection ($\mathrm{P_{d}}$), mean intersection-over-union (mIoU), and area under curve (AUC) for accuracy-based evaluation. In addition, we also use the frames per second (FPS), giga floating-point operations per second (GFLOPS), and parameter size as our efficiency-based evaluation.
IV-B Ablation Study in SISTD
IV-B1 Impact of Orthogonality
We conducted an ablation study to show the effects of using orthogonal versus non-orthogonal basis features, as summarized in Table I. The orthogonal basis features were obtained by applying singular value decomposition (SVD) to the original features. Experiments reveal that orthogonal features do not lead to any improvement in accuracy within our modules. On the contrary, they introduce a significant computational overhead. Consequently, the orthogonalization is unnecessary for our features.
IV-B2 Impact of the Number of Channels
To optimize channel dimension, we conduct a study on channel configurations, as detailed in Table II. We scale channel configurations from [4, 8, 16, 32] to [32, 64, 128, 256], yielding consistent improvements in mIoU. However, this comes at an increasing computational cost. The [8, 16, 32, 64] configuration achieves the optimal balance of performance across all metrics, and we adopt it as our channel configuration.
IV-B3 Module-wise Metrics Performance
We conduct ablation studies to validate the effectiveness of each module in our framework. The experimental results are summarized in Table III. All experiments are conducted on two SISTD datasets: NUDT-SIRST and IRSTD-1K. Using only the SD2M module achieves competitive results on NUDT-SIRST, confirming its effectiveness for target feature extraction. Additionally, the InceptionPooling module and SD3M module also enhance the network’s performance. Integrating all three modules achieves the optimal balance of performance.
TABLE IV: Quantitative comparison with different decomposition approaches on NUDT-SIRST and IRSTD-1K.
[…]
IV-C Comparison with State-of-the-art (SOTA) Methods in SISTD
TABLE V: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.