### v2

# Difference Decomposition Networks for Infrared Small Target Detection

Abstract
Weak target signal response is one of the key challenges in infrared small target detection (IRSTD), leading to missing detections, as some dim targets can be completely overwhelmed by the background. Recent studies have widely adopted difference-based methods, which utilize difference features in spatial and temporal domains. These methods rely on fixed and predetermined coefficients to combine these difference features and prevent the targets from overwhelmed by complex background, but fail to extract the correlation information between difference features and original information. To address this, we propose the Basis Decomposition Module (BDM) that can extract correlation information. Building on BDM, we develop the several variants for IRSTD and name them Difference Decomposition Modules (DDMs), which treat the normalized difference results as the basis features, combine them with the coefficient of these features projected onto the input, and achieve precise detection of small infrared targets without requiring a large number of parameters. Besides target enhancement from difference operations, DDMs also demonstrate robustness to diverse transformations. By integrating DDMs into U-shaped architectures, we develop two networks: the Spatial Difference Network (SD2Net) and the Spatiotemporal Difference Network (STD2Net) for single-frame IRSTD (SIRSTD) and multi-frame IRSTD (MIRSTD), respectively. With DDMs, they demonstrate performance competitive with most established networks on the SIRSTD and MIRSTD tasks. Our codes will be made public at: https://github.com/greekinRoma/IRSTD_HC_Platform.
I Introduction
Unlike visible light imaging, infrared sensing maintains reliable performance in adverse weather conditions, offering high detection reliability, concealment, and mobility. As one of the primary applications of infrared sensing, Infrared Search and Track (IRST) has demonstrated strong robustness under varying circumstances and serves as the foundation for several applications [Zhang2024IRPruneDet, Zhu2025PFF, iMoPKL2025, MICPL2025, yuan2025ASCNet]. Infrared small target detection (IRSTD) is a crucial task in IRST systems, garnering growing research interest. Nevertheless, the dim targets and complex backgrounds render IRSTD significantly more difficult than many other detection tasks [Zhu2024GSTUnet]. Therefore, IRSTD is a fundamental and challenging computer vision task, and merits further investigation [dai2025SeRankDet, zhang2025daaf, zhang2025coupled, yuan2025EDGSP].
In the past years, numerous model-based approaches have been proposed for single-frame IRSTD (SIRSTD) [kong2022LogTFNN, cao2021infrared, dai2017reweighted, wang2019miss]. Thanks to advances in deep learning, data-driven methods have become dominant in IRSTD, outperforming model-driven approaches [Zhang2025IRmamba, li2025hstnet, li2025mmidnet, zhu2025shifting]. However, spatial-based methods perform poorly when targets exhibit substantial visual similarity to their surroundings. In such circumstances, the primary difference between targets and false alarms is their motion. Thus, it is essential to introduce the motion information in complex backgrounds.
((a))
((b))
Figure 1: The standard CDC and frame-difference operations, both employing predetermined static coefficients to combine difference results. (a) We employ difference kernels to obtain difference features across various directions and combine them with static weights. (b) We extract motion features by subtracting the key frame from the reference frames and fuse them with static coefficients.
There are some general video object detection (GVOD) methods classifying targets from their false alarms [Deng2019Object, zhu2017deep, wu2019sequence, zhou2022transvod, wang2022ptseformer]. The infrared small targets are usually dim and move slowly in infrared streams [Duan2024Triple, zhang2025fsatfusion, zhang2025wife], but the GVOD methods could not highlight the motion information of such targets. Therefore, GVOD approaches are not suitable for IRSTD, motivating the exploration of new algorithms designed for infrared small targets. Currently, the standard technique for obtaining motion information for infrared small targets relies on the difference between key frames and reference frames [Li2025dtum].
In recent years, difference-based methods, such as central difference convolutions (CDCs) [Wu2025Salm] and difference-frame methods [Yan2023STDMANet], have gained an increasing attention. In these modules, detailed information is highlighted using high-pass kernels in either spatial or temporal dimensions [Zhang2021ECA]. However, both CDC-based approaches and other difference-based methods face a fundamental limitation in static coefficients, as illustrated in Figure 1.
From the perspective of basis decomposition, an input feature vector can be represented as a linear combination of a set of basis vectors, or basis features, with their corresponding coefficients. These coefficients are computed by projecting the basis features onto the input, thereby encoding the correlation between each basis feature and the input itself. We can exploit these coefficients to reconstruct the input, when the basis features are enough. If certain basis features are not enough, the input may lose the information those missed features represent,and highlight the retained ones.
Building on this idea, we design a unified and lightweight module for IRSTD, referred to as the Basis Decomposition Module (BDM). BDM selects specific features as basis features based on the given task, performs basis decomposition on the input using these features to obtain projection coefficients, and generates a feature as a combination of the weights and features. In this way, we could utilize different features as basis features for varied tasks.
For IRSTD, we modify the BDM and develop the Difference Decomposition Modules (DDMs) by utilizing the normalized difference results as our basis features. The output of each DDM is a combination of difference features, and thus strengthen the dim targets. Additionally, this design has two key properties that could enhance the robustness of our module:
  • (1)
    The normalization and difference process removes absolute intensity information, ensuring that the features only retain relative information. In this way, the module focuses on relative changes, unaffected by variations in the overall intensity of the input.
  • (2)
    The weights are dynamic and adjust based on the input. For example, when the input rotates in the spatial domain, the weights change accordingly. This design is particularly effective in space-based scenarios, such as satellite imagery, where perspective rotations are common. Therefore, this design makes it more flexible when dealing with spatial and temporal variations.
We introduce DDMs to a network and propose the Spatial Difference Network (SD2Net) and the Spatiotemporal Difference Network (STD2Net) for SIRSTD and MIRSTD, achieving a robust performance in their corresponding tasks using a limited number of parameters. To sum up, the contributions of this work are summarized as follows:
  1. 1.
    We introduce the BDM to enhance task-relevant features and suppress irrelevant information through basis decomposition, and thereby propose the DDMs for IRSTD.
  2. 2.
    We develop the SD2Net for SIRSTD to enhance high-frequency target features and suppress low-frequency clutter with a minimal number of parameters.
  3. 3.
    We further build the STD2Net, which is based on the SD2Net and effectively highlights moving targets while suppressing the static clutter for MIRSTD.
II Related Work
II-A Single-frame Infrared Small Target Detection
Existing SIRSTD methods can be categorized into two methodological paradigms: model-driven methods and data-driven methods.
Model-driven methods leverage domain-specific priors and bifurcate into three branches. Filter-based methods [Qin2019Facetkernel] enhance target features and suppress clutter with filters. Although these techniques perform well in simple scenes, they are susceptible to variations in window size and can easily fail when clutter exhibits similar textures or brightness. Human visual system-based methods [WeiYouLi2016Multiscale, Han2021WLCM] are based on visual characteristics of the visual system. However, these methods rely on handcrafted contrast metrics and struggle to generalize across diverse imaging conditions. Optimization-based methods [gao2013infrared, ZhangPeng2019Infrared] formulate IRSTD as an optimization problem to detect the targets by separating them from the background. Despite their interpretability, these models typically require iterative solvers, resulting in significant computation and limited scalability. Although these methods are computationally efficient and interpretable, their performance is often sensitive to parameter selection and unstable in complex, cluttered scenes.
Data-driven methods have addressed these issues of SIRSTD by learning discriminative features from large datasets. Dai et al. have successively proposed the Asymmetric Contextual Modulation (ACM) [DaiWuZhouBarnard2021Asymmetric] and Attentional Local Contrast Network (ALCNet) [dai2021attentional] for IRSTD. Dense Nested Attention Network (DNANet) [RenLiHanShu2021DNANet] supports adaptive interactions between feature layers to prevent the loss of targets. Furthermore, UIUNet [WuHongChanussot2023UIU-Net] employs a simple and effective “U-Net in U-Net” framework to enhance both global and local contrast information. Receptive-field and Direction-induced Attention Network (RDIAN) [SunBaiYangBai2023Receptive-Field] utilizes the characteristics of targets to solve the imbalance between targets and background. Infrared Small-target Detection U-Net (ISTDU-Net) [HouZhangTanXiZhengLi2022ISTDU-Net] introduces a fully connected layer in the skip connection to suppress the backgrounds with similar structures from the global receptive field. The Attention-guided Pyramid Context Network (AGPCNet) [ZhangCaoPuPeng2021AGPCNet] employs an attention-guided context block, providing the model with a perspective on both inner and global patches. The Spatial-Channel Cross Transformer Network (SCTransNet) [Yuan2024SCtransNet] utilizes the Transformer to extract practical global information.
Whereas previous models focused primarily on precision, our approach places greater emphasis on lightweight network design, aiming to strike a balance between computational cost and performance. This enables the single-frame model to remain sufficiently lightweight, creating favorable conditions for incorporating temporal information.
II-B Multi-frame Infrared Small Target Detection
Multi-frame frameworks can be categorized by their strategies for extracting temporal information.
Pre-extraction frameworks extract motion features before a SIRSTD network. Energy Accumulation (IFEA) [Du2022IFEA] and Spatiotemporal Differential Multiscale Attention Network (STDMANet) [Yan2023STDMANet] enhance the energy of targets by computing the difference between the current frame and reference frames before input into networks.
Spatiotemporal feature simultaneous extraction frameworks obtain the temporal and spatial information using modules in a single network. Sliced Spatiotemporal Network (SSTNet) [Chen2024SSTNet] utilizes the ConvLSTM node to extract the motion information of the small infrared targets. The Lightweight Asymmetric Spatial Feature Network (LASNet) [Chen2024LASNet] utilizes motion feature extraction and motion-affinity fusion for extracting temporal information. The Triple-domain Strategy (Tridos) [Duan2024Triple] employs frequency-aware enhancement to enhance the detection of dim infrared targets. The Local Motion Aware Transformer (LMAFormer) [huang2024lmaformer] utilizes local motion-aware attention to enhance the weak target motion information in MIRSTD. These end-to-end architectures can effectively integrate spatial and temporal cues; however, they typically require heavy computation and high memory usage, which limits their deployment in real-time infrared tracking systems. Furthermore, their focus on global aggregation often overlooks feature disentanglement, making it challenging to separate small targets from cluttered backgrounds.
Result-based extraction frameworks extract temporal information based on the results of single-frame neural networks. Direction-code Temporal U-shape Module (DTUM) [Li2025dtum] designs a new algorithm to extract the motion features from the results of SIRSTD networks. These approaches emphasize explicit motion cues and are straightforward to implement. Still, their reliance on handcrafted frame-difference operators makes them sensitive to background dynamics and sensor noise, often resulting in unstable motion estimation. The Recurrent Feature Refinement (RFR) framework [ying2025rfr] effectively captures long-term temporal dependencies based on the outputs of SIRSTD networks.
We adopt the result-based extraction framework as our foundation because its decoupling from feature-level fusion makes it inherently lightweight and easy to integrate into existing single-frame models.
((a))
((b))
Figure 2: (a) The procedure of basis decomposition, where $\vec{b_{1}}$, $\vec{b_{2}}$, and $\vec{b_{3}}$ are the basis vectors. In our study, we regard the results obtained by difference kernels or frame-difference methods as our basis vectors. (b) The illustration of the proposed module for signal recovery and task-relevant feature enhancement. First, we collect both task-relevant and task-irrelevant basis features into a basis pool according to the specific tasks. Then, the input is processed in two parallel branches for recovery and enhancement. The upper branch indicates the recovery process, where the input is reconstructed through basis decomposition. The lower branch outlines the enhancement procedure, which selects the task-relevant basis features to reconstruct the input.
II-C Difference-based Methods
In recent years, there have been increasingly difference-based methods in various methods [Su2025Rapid] to strengthen the detailed information in spatial and temporal domains.
In the spatial domain, researchers often use 2D convolution to extract spatial information. However, standard convolution is not sufficiently effective in obtaining the detailed information crucial for IRSTD. Therefore, there are some methods based on the classical descriptors, like Local Binary Pattern (LBP) [Ahonen2006LBP] and Gabor filters[JAIN1991Gaborfilter], to strengthen details [Xu2017LBPConv, luan2018GraborConv]. Building on this, central difference convolutions (CDCs) are proposed [yu2020CDC]. Chen et al. [Chen2024DEANet] parallel the vanilla convolutions and difference convolutions to obtain various features. Kim et al. [Kim2021Deformable] propose the Deformable Kernel Network, with adaptive neighbors and their corresponding weights.
In the temporal domain, some works choose to use the difference results between frames to obtain motion information[Yan2023STDMANet]. Besides, Geng et al. [GENG20253DCDC] propose the 3D central difference separable convolution and introduce the motion details into the difference convolution.
All of the studies demonstrated the strong power on highlighting the details and obtaining the motion information. Therefore, our method is justified.
III Methodology
III-A Preliminaries
III-A1 The procedure of basis decomposition
A vector $\vec{x}\in\mathbb{R}^{c_{\text{ful}}}$ can always be expressed as a linear combination of $c_{\text{ful}}$ orthonormal basis features.
As shown in the upper part of Figure 2(b), a vector $\vec{x}\in\mathbb{R}^{c_{\text{ful}}}$ can be recovered by $c_{ful}$ orthonormal vectors:
$c_{ful}=dim(\vec{x}),$ (1)
$s_{i}=\vec{b}_{i}^{\top}\vec{x}\,$, (2)
$\vec{x}=\sum_{i=1}^{c_{\text{ful}}}s_{i}\vec{b}_{i}\,$, (3)
where $s_{i}$ denotes the coefficient of $\vec{x}$ projected onto the basis vector $\vec{b}_{i}$. $\top$} is the transposition operation, and $c_{\text{ful}}$ is the dimension of the input vector. Additionally, dimdim denotes the dimensionality of the vector $\vec{x}$.
((a))
((b))
Figure 3: Robustness of our methods. (a) Illustration of robustness to data value variations. The representation is constructed based on normalized relative differences between neighboring elements, making it insensitive to uniform shifts or scale changes in the input values. (b) Illustration of rotation invariance. Local patterns under different input directions are aligned to an orientation, resulting in identical weight arrangements for rotated versions of the same pattern.
Furthermore, the result can be constructed solely from the selected basis elements and the irrelevant basis features are removed, thereby enhancing the task-relevant features, as illustrated in the lower portion of Figure 2(b). In this way, we can strengthen the important features by reducing interference from non-essential information.
Therefore, for a given task, it is not necessary to retain all basis features, and we could select the task-relevant basis vectors to obtain $\vec{x}$ and approximate the input vector $\vec{x}$:
$\vec{y}=\sum_{i=1}^{c_{\text{tas}}}s_{i}\vec{b}_{i}\quad\text{where}\quad s_{i}=\vec{b}_{i}^{\top}\vec{x}\,$, (4)
where $c_{\text{tas}}$ is the number of the task-relevant basis features.
III-A2 Relaxing Constraints:
Basis vectors must be both normalized and orthogonal. As shown in Table I, our method does not require the basis features to be strictly orthogonal, because forcing $\{\vec{b}_{k}\}$ to be orthogonal would require an additional decorrelation step, which would mix these simple local difference patterns into more global filters and destroy their physical interpretation. By contrast, normalization plays essential roles, which removes the scale ambiguity between a basis vector and its coefficient, prevents basis features with larger norms from dominating the decomposition, and improves robustness. Thus, the orthogonality constraints are relaxed, while the normalization requirement is retained.
Figure 4: Overview of the BDM.
III-A3 Difference Decomposition
When it comes to IRSTD, where detection relies more on local contrast than on absolute intensity, it is natural to construct features from difference information. Consequently, we treat the difference features, obtained by difference kernels or frame difference methods, as basis feature vectors and combine them to form $\vec{y}$ and approximate $\vec{x}$. Ultimately, $\vec{y}$, the fusion of difference features, enhances the high-frequency components that contain information about infrared small targets. In our module, the decomposition coefficients are not learned as fixed parameters. Instead, they are computed dynamically from the current input. Thus, the coefficients adapt to each sample and each location, rather than being shared globally across all inputs. Difference decomposition has two main properties:
  • 1.
    We begin by performing difference extraction and normalization to obtain the basis vectors from the relative differences, as illustrated in Figure 3(a). This approach to difference decomposition mitigates the impact of overall intensity changes in features, as each basis is defined based on differences rather than absolute values.
  • 2.
    The output is formed by combining the basis features with their corresponding coefficients, where the coefficients are computed from the input features and the basis feature group. Consequently, if both the input features and the basis feature group remain unchanged, the output of our module will also remain unchanged. For example, under spatial difference decomposition, even when the input undergoes a rotation when the basis features remain fixed, the output retains the origin result, as shown in Figure 3(b).
Therefore, due to the difference decomposition method, it offers more robust and effective performance compared to conventional techniques, such as CDCs and frame-difference methods.
III-B Basis Decomposition Module
III-B1 Overall Procedure
We propose the BDM based on the above discussion, as shown in Figure 4.
((a))
((b))
Figure 5: Detailed structure of the SD2M and SD3M used in Figures 7 and 9 respectively, are constructed based on the BDM. SD2M is utilized for the feature extraction without
Figure 6: Our backbone adopts a three-stage U-Module structure, where each U-Module is an encoder–decoder architecture. The first U-Module extracts coarse information, the second U-Module performs as the main feature-extraction stage to obtain the refined information, and the third U-Module conducts fine-grained enhancement.
Figure 7: The overall architecture of SD2Net, which incorporates the SD2M and SD3M modules.
Firstly, we employ the Feature Extraction Module (FEM) and Basis Feature Extraction Module (BFEM) to obtain the original feature $T\in\mathbb{R}^{B\times G\times 1\times F}$ and the basis features $P_{i}\in\mathbb{R}^{B\times G\times 1\times F}$ from $I\in\mathbb{R}^{B\times G\times 1\times F_{inp}}$:
$T$ $=\mathrm{FEM}(I)\,$, (5)
$P_{1},\dots,P_{c}$ $=\mathrm{L2Norm}(\mathrm{BFEM}(I))\,$, (6)
where $B$ is the batch size, $F$ is the feature dimension, and $F_{inp}$ is the input feature dimension. $G$ represents the number of feature groups, as we can regard several features across channels as one group. $\mathrm{L2Norm}$ refers to L2 normalization to ensure each basis feature has the same influence on $\vec{y}$.
Then, We concatenate the $\{P_{i}\}$ to obtain $P\in\mathbb{R}^{B\times G\times c\times F}$:
$P=[P_{1};\dots;P_{c}]\,$, (7)
where $c$ refers to the number of input channels, while $\cdot$$\cdot$] is the concatenation operator applied along the certain dimension.
Figure 8: Architecture of the DFEM, which employs eight difference convolution kernels with $n$ dilation rates to extract multi-scale basis features.
After that, we utilize the $T$ and $P$ to obtain the tensor $S\in\mathbb{R}^{B\times G\times 1\times c}$ from concatenating decomposition coefficients $S_{i}\in\mathbb{R}^{B\times G\times 1\times 1}$ for different basis features $P_{i}$:
[…]
where $[,]$ concatenates elements along the last dimension.
Finally, the weight $S_{i}$ is assigned to each basis feature $P_{i}$:
$\begin{split}O=S\cdot P=(T\cdot P^{\top})\cdot P=\sum_{i=1}^{c}(T\cdot P_{i}^{\top})\cdot P_{i}\,.\end{split}$ (9)
In short, FEM extracts features containing both salient and redundant information, and BDM approximates the FEM results by linearly combining the basis features from BFEM. This approximation suffers from information loss. Therefore, we must retain the essential information and disregard the redundant information. As the definition of essential information is task-dependent, BFEM requires corresponding adjustments for specific tasks.
In processing two-dimensional feature maps, spatial dimensions (height $H$ and width $W$) are flattened into tokens. For an input $\mathbf{I}\in\mathbb{R}^{B\times C\times H\times W}$, this yields basis features of size $B\times C\times c\times(HW)$, where $HW$ and $C$ map to dimensions $F$ and $G$.
III-B2 Difference Decomposition Module
The primary difference between the proposed DDMs and the original module lies in the use of the BFEM. This module dynamically derives difference features across spatial and temporal dimensions using a variety of techniques, including central difference convolutions and the difference between two frames.
III-C Spatial Difference Decomposition Network
As illustrated in Figures 5(a) and 5(b), we adjust the BDM for IRSTD and introduce two DDMs, named Spatial Difference Decomposition Module (SD2M) and Spatial Difference Decomposition Downsampling Module (SD3M), to enhance the targets and suppress clutter. In addition, we add those modules to our backbone, as shown in Figure 6, and propose the SD2Net, as illustrated in 7.
SD2M serves as a difference information extractor that maintains resolution. SD3M acts as a detail-preserving downsampling module, embedded into a multi-branch module.
III-C1 Backbone
As illustrated in Figure 6, we design an architecture composed of three U-Modules. Given an input $I$, the first U-Module produces a coarse representation $f_{1}=\mathcal{U}_{1}(I)$. The second U-Module extract features further and works as the main feature extraction layer: $f_{2}=\mathcal{U}_{2}(f_{1})$. The third U-Module works for final result enhancement: $O=\mathcal{U}_{3}(f_{2})$. The overall mapping is
$O=\mathcal{U}_{3}\big(\mathcal{U}_{2}(\mathcal{U}_{1}(I))\big).$ (10)
III-C2 Spatial Difference Decomposition Module
As depicted in Figure 5(a), we employ a $1\times 1$ convolution layer followed by a reshape operation ($\mathrm{Reshape}$) as FEM to extract the original feature $T_{m}\in\mathbb{R}^{B\times C\times 1\times(HW)}$:
$T_{m}=\mathrm{Reshape}(\mathrm{Conv}_{1\times 1}(I))\,$, (11)
where H, W, and C denote the height, width, and channel dimension. The Difference Feature Extraction Module (DFEM) utilizes difference kernels to extract high-frequency basis features ${P^{d_{j}}_{i}}\in\mathbb{R}^{B\times C\times 1\times(HW)}$ (where i and j index the kernel type and dilation rate, respectively). These features are then used to form the tensor $P_{m}\in\mathbb{R}^{B\times C\times(8n)\times(HW)}$:
$P_{m}$ $=\mathrm{L2Norm}(DFEM(I))\,$, (12)
$DFEM(I)$ $=[P^{d_{1}}_{1};\ldots;P^{d_{n}}_{1};\ldots;P^{d_{n}}_{8}]\,$, (13)
$P^{d_{j}}_{i}$ $=\text{Reshape}(\mathrm{Conv}_{K_{i}}^{d_{j}}(I))\,$, (14)
where $\mathrm{Conv}_{K_{i}}^{d_{j}}$ is the difference convolution with kernels $K_{i}$ ($i\in\{1,\dots,8\}$), as shown in Figure 8, and dilation rates $d_{j}$ ($j\in\{1,\dots,n\}$) to extract the high-frequency features as our basis features.
We utilize $P_{m}$ and $T_{m}$ to obtain $O_{m}\in\mathbb{R}^{B\times C\times H\times W}$ according to Equation 9:
$O_{m}=\mathrm{Reshape}((T_{m}\cdot P_{m}^{\top})\cdot P_{m}\,)\,$, (15)
where $O_{m}$ is a combination of difference features. In the process, the background information is suppressed by overlooking the low-frequency basis features.
Figure 9: The architecture of Inception Pooling.
Figure 10: The structure of STD2Net comprises two main components: the TFEM is employed to extract difference basis features in the temporal domain. At the same time, the FEM is utilized to fuse the reference frames with the key frame to obtain the features with temporal information.
III-C3 Inception Pooling.
While SD3M could keep detailed information, the downsampling process itself requires broader contextual information beyond these details. Therefore, we utilize multiple branches to maintain various information.
To maintain feature diversity at various levels, Inception Pooling splits the feature channels into groups and applies distinct operations to each group, as shown in Figure 9.
First, we split the input $I$ along the channel dimension into three equal-sized tensors $C_{t}\in\mathbb{R}^{B\times(C/3)\times H\times W}$.
$C_{1},C_{2},C_{3}=\mathrm{Split}(I)\,$, (16)
After that, we use MaxPooling $\mathrm{Max}_{\text{2\texttimes 2}}$, AvgPooling $\mathrm{Avg}_{\text{2\texttimes 2}}$, and $\text{SD}^{\mathrm{3}}\text{M}$ to downsample the input and concatenate the results in channel dimension by concatenation ($\mathrm{CAT}$).
$O=\text{CAT}(\mathrm{Max}_{\text{2\texttimes 2}}(C_{1}),\mathrm{Avg}_{\text{2\texttimes 2}}(C_{2}),\text{SD}^{\mathrm{3}}\text{M}(C_{3}))\,$, (17)
III-C4 Spatial Difference Decomposition Downsampling Module.
As illustrated in Figure 5(b), the SD3M shares the similar architecture with SD2M, differing in the original feature $T_{d}\in\mathbb{R}^{B\times C\times 1\times(HW/4)}$ and the basis features $P_{d}\in\mathbb{R}^{B\times C\times 4\times(HW/4)}$, which is computed as follows:
$T_{d}$ $=\mathrm{Reshape}(\mathrm{Conv}_{\mathrm{1\times 1}}(\mathrm{Avg}_{\mathrm{2\times 2}}(I)))\,$, (18)
$P_{d}$ $=\mathrm{L2Norm}(\text{DFED}(I))\,$, (19)
where DFED denotes the Difference Feature Extraction Downsampling, formulated as follows:
$\text{DFED}(I)$ $=[P_{max};P_{D_{1}};P_{D_{2}};P_{D_{3}}]\,$, (20)
$P_{max}$ $=\mathrm{Reshape}(\mathrm{Max}_{\mathrm{2\times 2}}(I))\,$, (21)
$P_{D_{k}}$ $=\mathrm{Reshape}(\mathrm{Conv}^{1}_{D_{k}}(I))\,$, (22)
where $\mathrm{Conv}^{1}_{D_{k}}$ denotes a convolution with stride 2, dilation ratio of 1, and the kernel $D_{k}$, which is defined as: $D_{1}=[1,1;-1,-1],D_{2}=[1,-1;1,-1],D_{3}=[1,-1;-1,1]$.
At last, we obtain the output of $O_{d}\in\mathbb{R}^{B\times C\times H/2\times W/2}$:
$O_{d}=\mathrm{Reshape}((T_{d}\cdot P_{d}^{\top})\cdot P_{d}\,)\,.$ (23)
where $O_{d}$ is a linear sum of high-frequency difference features and MaxPooling results from I.
III-D Spatiotemporal Difference Decomposition Network
Since background clutter exhibits random motion patterns, while targets move consistently. Therefore, we introduce TD2M to motion information and propose STD2Net, as shown in Figure 10. The basis features extracted by Temporal Difference Decomposition Module (TD2M) consist of both the differences in SD2Net’s results between the current frame and the reference frames, and the current frame’s SD2Net output. The former captures motion information, while the latter provides current state information.
Figure 11: Visual results from varied data-driven methods. Green, red, and yellow boxes indicate detected targets, missed targets, and false alarms.
At first, we extract the input features $I_{k}\in\mathbb{R}^{B\times C\times H\times W}$ (k = 1,…,$N$) for each of the $N$ consecutive frames using SD2Net, and concatenate them in channel dimension to obtain $I\in\mathbb{R}^{B\times(NC)\times H\times W}$ :
$I=\mathrm{CAT}(I_{1},\dots,I_{N})\,$, (24)
where the last frame is the current frame and the preceding N-1 frames are reference frames.
Then, we utilize the TD2M to extract the relationships across these consecutive frames. We obtain original features $T_{t}\in\mathbb{R}^{B\times C\times 1\times(HW)}$:
$T_{t}=\mathrm{Reshape}(\mathrm{Conv}_{1\times 1}(I))\,.$ (25)
After that, we employ the Temporal Feature Extraction Module (TFEM) and the normalization layer to obtain the basis features $P_{t}\in\mathbb{R}^{B\times C\times N\times(HW)}$:
$P_{t}$ $=\mathrm{L2Norm}(\mathrm{TFEM}(I))\,$, (26)
$\mathrm{TFEM}(I)$ $=[P_{D_{1}};..;P_{D_{N-1}};P_{C}]\,$, (27)
$P_{D_{i}}$ $=\mathrm{Reshape}(I_{N}-I_{i})\,$, (28)
$P_{C}$ $=\mathrm{Reshape}(I_{N}),$ (29)
where $P_{D_{i}}$ denotes the difference between the current frame $I_{N}$ and the reference frame $I_{i}$. $P_{C}$ is the state of the current frame.
Later, we approximate $T_{t}$ with a combination of $P_{t}$ and obtain the result $O_{t}\in\mathbb{R}^{B\times C\times H\times W}$:
$O_{t}=\mathrm{Reshape}((T_{t}\cdot P_{t}^{\top})\cdot P_{t})\,.$ (30)
At last, we use a U-Module in Figure 7 to further extract information and a convolution layer to adjust the output channel dimension.
Although the dot-product form resembles attention, BDM differs fundamentally: it performs linear basis reconstruction rather than parametric attention weighting. The coefficients arise from geometric alignment between features and fixed bases, ensuring interpretability and low computational cost.
III-E Loss Function
We employ the soft Intersection over Union (IoU) loss for training our model, which is defined as:
$\mathcal{L}=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (31)
where $g_{i,j}\in[0,1]$ and $p_{i,j}\in[0,1]$ represent the ground truth and predicted probability after Sigmoid at $(i,j)$.
IV Experiments
We present the experiment details, validate modules via ablation studies, and demonstrate the performance of BDM’s adaptations through comparisons between SIRSTD and MIRSTD.
IV-A Experimental Details
IV-A1 Experiment Dataset
For the task of the SIRSTD, we employ two widely recognized public benchmarks: NUDT-SIRST [RenLiHanShu2021DNANet] and IRSTD-1K [Zhang2022ISNet]. To evaluate performance on the more complex task of MIRSTD, we benchmark our network against existing SOTA methods using the comprehensive NUDT-MIRSDT dataset [Li2025dtum]. To ensure a fair and consistent evaluation across all experiments, we divide each dataset into two parts: 80% of the images are randomly selected to constitute the training set, while the remaining 20% are reserved as the test set for the final performance assessment. All comparative methods were evaluated on the same dataset to ensure a fair comparison.
TABLE I: Impact of orthogonality and normalization.
[…]
TABLE II: Impact of SD2Net’s dilation configs on NUDT-SIRST dataset. We bold the best results and underline the second best.
[…]
TABLE III: Impact of SD2Net’s channel configs on NUDT-SIRST dataset. We bold the best results and underline the second best.
[…]
TABLE IV: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.
[…]
IV-A2 Implementation Details
The loss function for all models is the IoU loss [yu2016unitbox]. All experiments are conducted on Ubuntu 20.04 LTS. We train SIRSTD models on an NVIDIA GeForce RTX 4080 GPU for 400 epochs, using the Adam optimizer with a learning rate of $5\times 10^{-4}$. The batch size is set to 4. The learning rate is decayed to $5\times 10^{-5}$ at epoch 200 and further reduced to $5\times 10^{-6}$. Our MIRSTD architecture runs on two 24GB GeForce RTX 3090 GPUs. We train for 20 epochs with the Adam optimizer and an initial learning rate of 0.001. Furthermore, we set the temporal window size $N$ of STD2Net to 5. We run each algorithm three times and use the average outcome as the final result. All models, including the baselines, are trained under our experimental configuration to ensure a fair comparison.
IV-A3 Evaluation Metrics
We adopt four evaluation metrics: false alarm rate ($\mathrm{P_{a}}$), probability of detection ($\mathrm{P_{d}}$), mean intersection-over-union (mIoU), and area under curve (AUC) for accuracy-based evaluation. In addition, we also use frames per second (FPS), giga floating-point operations per second (GFLOPS), and parameter size as our efficiency-based evaluation metrics.
IV-B Ablation Study in SIRSTD
IV-B1 Impact of Orthogonality and Normalization
We conducted an ablation study to show the effects of using orthogonal versus non-orthogonal basis features, as summarized in Table I. The orthogonal basis features were obtained by applying singular value decomposition (SVD) to the original features. Experiments reveal that orthogonal features do not lead to any improvement in accuracy within our modules. On the contrary, they introduce a significant computational overhead. However, the normalization is the key part of our method. Consequently, we retain the normalization and relax the orthogonality constraint.
IV-B2 Impact of the Dilation Ratios
To determine the optimal setting, we performed an ablation study on dilation ratios, systematically evaluating the performance of various configurations, as shown in Table II. The experiment results indicated that the [1, 3] configuration significantly outperformed the others. Based on these findings, we adopted [1, 3] as our dilation configuration. We apply a dilation ratio of 3 to larger targets and a ratio of 1 to smaller targets.
IV-B3 Impact of the Number of Channels
To optimize channel dimension, we conduct a study on channel configurations, as detailed in Table III. We scale channel configurations from [4, 8, 16, 32] to [32, 64, 128, 256], yielding consistent improvements in mIoU. However, this comes at an increasing computational cost. The [8, 16, 32, 64] configuration achieves the optimal balance of performance across all metrics, and we adopt it as our channel configuration.
IV-B4 Module-wise Metrics Performance
The SD2M works for enhancing the high-frequency targets. Additionally, the SD3M and InceptionPool are utilized for downsampling, which removes less task-relevant information. We conduct ablation studies to validate the effectiveness of each module in our framework. The experimental results are summarized in Table IV. All experiments are conducted on two SIRSTD datasets: NUDT-SIRST and IRSTD-1K. Using only the SD2M module achieves competitive results on NUDT-SIRST, confirming its effectiveness for target feature extraction. Additionally, the InceptionPooling module and SD3M module also enhance the network’s performance. Integrating all three modules achieves the optimal balance of performance. Therefore, all modules are effective.
TABLE V: Quantitative comparison with different decomposition approaches on NUDT-SIRST and IRSTD-1K.
[…]
IV-C Comparison with State-of-the-art (SOTA) Methods in SIRSTD
TABLE VI: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.